Kyky232:

What is the range of g(x)=3^(-x)

Hashtag:

Do you know how to find a range?

1 attachment
Kyky232:

no

DRAK0:

Remove parentheses. 3 ^-x Use Negative Power Rule X^-6=1/x^a. ​ 1/3^x

surjithayer:

\[g(x)=3^{-x}=\frac{ 1 }{ 3^{x} }\ge 0 , ~\forall ~values~of~x\] Range:- all positive real numbers.